OpenStudy (anonymous):

Find the slope of the line that contains (9, -1) and (9, 9)

OpenStudy (saifoo.khan):

(9+1)/(9-9) equals zero.

OpenStudy (saifoo.khan):

means the line is is parallel to x-axis!

OpenStudy (anonymous):

Using the slope formula: \[slope = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \frac{9+1}{9-9} = \frac{10}{0}\] Your slope is undefined (not zero)

OpenStudy (anonymous):

the line is vertical so its slope is UNDEFINED

OpenStudy (anonymous):

its parallel to the y axis, not x axis

OpenStudy (saifoo.khan):

Yeah, its parellel to y-axis/./ My bad.

OpenStudy (saifoo.khan):

its undefined.